# Statistical testing of random number generators and their improvement using randomness extraction

### v1

@@ Abstract
Random number generators (RNGs) are notoriously hard to build and test, especially in a cryptographic setting. Although one cannot conclusively determine the quality of an RNG by testing the statistical properties of its output alone, running numerical tests is both a powerful verification tool and the only universally applicable method. In this work, we present and make available a comprehensive statistical testing environment ($\mathsf{STE}$) that is based on existing statistical test suites. The $\mathsf{STE}$ can be parameterised to run lightweight (i.e. fast) all the way to intensive testing, which goes far beyond what is required by certification bodies. With it, we benchmark the statistical properties of several RNGs, comparing them against each other. We then present and implement a variety of post-processing methods, in the form of randomness extractors, which improve the RNG’s output quality under different sets of assumptions and analyse their impact through numerical testing with the $\mathsf{STE}$.
I Introduction
@@
The notion of randomness plays an important role in numerous fields, ranging from philosophy to science. In science, it is used in optimisation and numerical integration (e.g. using the Monte Carlo method), algorithm randomisation or cryptography. Although there is something universal about the concept of randomness, its definition varies strongly depending on the context in which it is used. In cryptography, for example, random numbers should be unpredictable after they have been generated, even by an adversary potentially possessing information about the random number generator (RNG) that the user does not have. Therefore, randomness – or unpredictability – from the perspective of the RNG user and from the perspective of a hypothetical adversary are fundamentally different. However, if the output of the RNG exhibits patterns that are detectable by the user, then these patterns also imply predictive power from the perspective of the adversary, since the adversary needs to be considered to have at least as much information as the user. In that sense, unpredictability from the user’s perspective is a necessary (but not sufficient) condition for the unpredictability of an adversary. This idea motivates the use of numerical testing of a RNG’s outputs, which serves as a means of randomness validation, i.e. to detect failure to generate randomness.
@@
Because numerical testing is a useful implementation check and the only universally applicable method to test different RNGs, it is an essential part of getting a cryptographic RNG certified by standards bodies, for example the National Institute of Standards and Technology (NIST) or the Bundesamt für Sicherheit in der Informationstechnik (BSI). The purpose of this certification process is for a third party, for example NIST, to provide assurance that the RNG has been built and tested according to the best practices. In fact, in almost all applications related to cryptography, most companies see the certification by NIST (or another equivalent body) of the RNG as a requirement for using it. NIST and BSI’s standards require both a detailed modelling of the underlying physical process and numerical testing of the RNG’s output statistics in order for a hardware RNG to be compliant. Partly because of this, numerous statistical test suites have been developed. The best known are NIST’s rukhin2001statistical and the Dieharder marsaglia2008marsaglia ; brown2018dieharder suite, but other useful ones exist e.g. l2007testu01 ; ent ; doty2018practrand . Together, the tests contained within these suites enable the comprehensive analysis of a wide array of statistical characteristics. Despite their usefulness, these suites are often complicated to use and their output subtle to analyse. Moreover, it is desirable to combine tests from different suites in order to push the statistical testing further.
@@
Because of this, in our work we select and parameterise specific statistical tests from existing suites to create a statistical testing environment ($\mathsf{STE}$) with a tunable intensity from lightweight to intense, with a recommended setting offering a good trade-off between its computational cost and its effectiveness in detecting statistical bias. Due to its ability to be tuned, the suite enables testing beyond that of the individual suites it selects from, and allows for more rigorous testing of an RNG compared to the requirements of standards bodies. We make our testing environment easy to use and openly available at https://github.com/CQCL/random_test. Using the $\mathsf{STE}$’s most intense version, we test the output of three different RNGs representative of those used across commercial applications, allowing us to benchmark them against each other. We also provide a framework for the analysis of the results of the overall numerical test results, which is not obvious otherwise.
@@
We then use the $\mathsf{STE}$ to test the impact of different techniques to improve the RNG’s output by post-processing it. For this, we present and implement a variety of randomness extraction methods, with the mathematical algorithms taken from Cryptomite, the software library of randomness extractors that we developed in cryptomite . Extraction methods are applied to the output of the RNG to improve its quality, e.g. by removing any bias or dependencies between bits. The set-up is illustrated in Figure 1. Each of the randomness extraction methods we present relies on its own set of assumptions, which can be compared against each other. Such assumptions are, for example, additional structure of the RNG’s outputting process, i.e. that each output bit is produced in an identical and independently distributed (I.I.D) manner, or the need for a short pre-existing (near-)perfect random bit string as a resource. These assumptions need to be justified in practice, and statistical testing is used to assess whether each post-processing method was successful from a statistical perspective.
@@
Figure 1: This figure illustrates our implementation set-up. The black box represents one of the initial RNGs we test and the dashed box denotes the new, in principle improved, RNG with additional post-processing applied.
I.1 Related work
@@
Statistical testing of RNGs has a long history, dating back to the implementation of the Diehard CD-ROM tests in the 1990s marsaglia2008marsaglia . Since then, two main research directions have emerged, which are both relevant to our work. First, researchers developed other test suites, such as the NIST Statistical Test Suite rukhin2001statistical , the TestU01 suite l2007testu01 , ENT ent , and PractRand doty2018practrand , all of which we utilise for this work. Second, empirical testing of specific RNGs was performed, such as soto1999statistical ; tsvetkov2011empirical , which analysed and compared the results of numerous statistical tests on a variety of pseudo-RNGs (PRNGs). Other works have considered empirical testing of so-called true-RNGs (TRNGs), for example hamburg2012analysis which tests the statistical properties of the entropy source in Intel’s Ivy Bridge TRNG jun1999intel and tsoi2007high ; zhang2017640 ; williams2010fast ; sun2018random ; cho2020random which develop, implement and statistically test different TRNGs. Recently, this has been extended to quantum RNGs (QRNGs), for example, xu2019high ; o2023operation ; jacak2021quantum ; keshavarzian20233 and hurley2020quantum , in which the authors’ extensively statistical test an ID Quantique QRNG quantique2004quantis . Other works give a universally applicable RNG statistical testing framework, such as crocetti2023review ; seyhan2022classification .
@@
Randomness extraction also has a rich literature, see shaltiel2011introduction for an introduction. In cryptographic randomness standards, e.g. in NIST’s SP 800 rukhin2001statistical , so-called conditioners are standardised, whose role is similar to randomness extractors [^1] Randomness extractors can be understood as conditioners that have information-theoretic security, i.e. do not rely on computational assumptions on the adversary.. These conditioners are therefore the only post-processing that has been vetted for use by governing bodies and are the most commonly used as a consequence. To the best of our knowledge, our work represents the first attempt at comparing the effect (from a statistical perspective) of different post-processing methods.
I.2 Summary of results
In summary, our results and observations are:
  • To make available our $\mathsf{STE}$, a powerful, flexible and easy to use statistical testing environment together with the framework to analyse its results.
@@ I.2 Summary of results
  • To intensely test the output statistics of three different RNGs: the 32-bit Linear Feedback Shift Register (LFSR) PRNG, Intel’s RDSEED TRNG, and IDQuantique’s ‘Quantis’ QRNG. We show the failure of two of them and provide evidence that one behaves well from a statistical perspective, extending and confirming the results of hurley2020quantum ; hamburg2012analysis .
@@
  • To present and implement a variety of post-processing techniques, in the form of randomness extractors, to improve the quality of the outputs for each of the three RNGs. This set of post-processing methods is made of four levels, requiring increasingly more sophisticated implementations: deterministic (level 1), seeded (level 2), two-source (level 3) and physical (level 4) extractors. It goes significantly further than the study and comparison of different types of extractors in kwok2011comparison and ma2013postprocessing , which respectively only focus on deterministic or seeded extractors. Together with the software library Cryptomite, presented separately in cryptomite and which we use for our extractor implementations, this allows RNG developers to carefully choose and implement suitable post-processing.
@@
  • To use the $\mathsf{STE}$ to intensely analyse the effect that each level of the post-processing has on the output of the different RNGs. Our main observations are that:
    • The RNGs that failed numerical testing without post-processing still fail when simpler post-processing methods (level 1) are applied, although an improvement is indeed observed.
@@
    • All our implementations at levels 2, 3 and 4 successfully post-process, from a statistical perspective, the output of the three RNGs that we used.
@@
    • Low entropy sources, for example the post-processed 32-bit LFSR, can successfully pass our intense statistical testing when the right post-processing is applied. This is indeed unsurprising based on the existence of PRNGs, but the poorness of the PRNG used illustrates the limitation of statistical testing when performed alone, i.e. without a precise model and justification for the unpredictability of the underlying physical process.
We note that there are sometimes important differences between the ideal version of a certain extraction method and its actual implementation. Because of this, we ensure to explicitly state both fundamental and added implementation assumptions at each level.
II Tools and Definitions
As we stated in the introduction, randomness is different when considering the perspective of a user or that of an adversary. The difference between statistical and cryptographic randomness can be understood by considering a hypothetical game in which an adversary tries to distinguish the real output of an RNG from that of an ideal RNG, i.e. one whose output distribution is uniformly distributed. The difference then lies, mainly, in the information available to the adversary in the game [^2] In practice, another difference is the computational power available, since a user can only run limited statistical testing when there is no such limitation for a potentially computationally unbounded adversary.. In the case of cryptographic randomness, the adversary has access to additional (aka side) information that is not in the hands of the user, i.e. its predictive power cannot be quantified directly by studying the output’s statistical properties alone. Examples of such side information include a better model of the underlying physical process (the entropy source) or of the device’s environment, or even information leaked through side-channels. For statistical randomness, one assumes that there is no additional side information, since only the statistical properties of the output is analysed.
As an example, pseudo-RNGs (PRNGs) are mathematical algorithms that expand an initial, short, string of random numbers into a larger output string. The initial random string is called the seed of the PRNG and needs to be obtained independently. The (larger) output of the PRNG is then indistinguishable from the uniform distribution to an adversary who is unable to solve a given computational problem efficiently, e.g. the learning with errors (LWE) problem. In this case, there is a big difference between an adversary with no additional information and one who knows the seed. In one case where the seed is unknown, the output is statistically indistinguishable from the uniform distribution (and therefore should pass statistical tests). In the other case, however, the output can be recomputed directly from the seed and therefore be distinguished from the ideal output. This is why it is crucial to keep the seed private when using a PRNG.
@@ II Tools and Definitions
In our work, we study the output distribution of different RNGs directly, without considering additional information i.e. we test for statistical randomness. We then study, in the same way, the effect of different post-processing methods applied to the output of the RNG. We consider the specific case of bits, i.e. the RNG outputs an $n$ length bit string $X\in\{0,1\}^{n}$, for simplicity - although one can also study RNGs whose output alphabet is larger. In some cases, we talk about sizes in bytes, where one byte is eight bits. We denote the random variable $X$ and its specific realisation $x$ for $X=x$. The set-up is illustrated in Figure 2.
@@
Figure 2: An illustration of the set-up we consider. An RNG generates a bit string $X=x$ of length $n$. In this work, we first study the statistical properties of the realisation $x$ of the (random variable) $X$. Then, we analyse the effect of different post-processing methods applied to it.
@@
The amount of randomness a random variable $X$ has is captured by its min-entropy $\mathrm{H}_{\infty}(X)$.
Definition 1 (Min-entropy).
The min-entropy, $k$, of a random variable, $X\in\{0,1\}^{n}$, is defined as
@@
This can be interpreted as the minimum amount of randomness, in bits, a variable $X$ has, when there is no side information available. The quantity $P_{g}=\max\limits_{x\in\{0,1\}^{n}}\Pr(X=x)$ is the guessing probability of $X$. Since RNGs output sequentially, we generalise this definition to consider the min-entropy of the current random variable conditioned on all previously produced random variables. This is known as block min-entropy.
Definition 2 (Block min-entropy).
A set of random variables $X_{i}\in\{0,1\}^{n}$ for $i\in\mathbb{N}$ is said to have block min-entropy [^3] This definition can easily be generalised to the case where each block, or random variable, $X_{i}$ has different size $n_{i}$, i.e. $X_{i}\in\{0,1\}^{n_{i}}$, and min-entropy $k_{i}$. $k$, if

### v2

# statistical testing; random number generation; randomness extractors; information-theoretic security

@@ Abstract
Random number generators (RNGs) are notoriously challenging to build and test, especially for cryptographic applications. While statistical tests cannot definitively guarantee an RNG’s output quality, they are a powerful verification tool and the only universally applicable testing method. In this work, we design, implement, and present various post-processing methods, using randomness extractors, to improve the RNG output quality and compare them through statistical testing. We begin by performing intensive tests on three RNGs—the 32-bit linear feedback shift register (LFSR), Intel’s ‘RDSEED,’ and IDQuantique’s ‘Quantis’—and compare their performance. Next, we apply the different post-processing methods to each RNG and conduct further intensive testing on the processed output. To facilitate this, we introduce a comprehensive statistical testing environment, based on existing test suites, that can be parametrised for lightweight (fast) to intensive testing.
keywords:
statistical testing; random number generation; randomness extractors; information-theoretic security
[…]
1 Introduction
@@
The notion of randomness plays an important role in numerous fields, ranging from philosophy to science. In science, it is used in optimisation and numerical integration (e.g., using the Monte Carlo method), algorithm randomisation, or cryptography. Although there is something universal about the concept of randomness, its definition varies substantially depending on the context in which it is used. In cryptography, for example, random numbers should be unpredictable, in the sense that they should be indistinguishable from uniformly distributed and secret ones, even by an adversary potentially possessing information about the random number generator (RNG) that the user does not have. Therefore, randomness—or unpredictability—from the perspective of the RNG user and from the perspective of a hypothetical adversary is fundamentally different. However, if the output of the RNG exhibits patterns that are detectable by the user, then these patterns also imply predictive power from the perspective of the adversary, since the adversary needs to be considered to have at least as much information as the user. In this sense, unpredictability from the user’s perspective is a necessary (but not sufficient) condition for the unpredictability of an adversary. This idea motivates the numerical testing of RNGs’ outputs, which serves as a means of randomness validation, i.e., to detect failures to generate randomness.
@@
Because numerical testing is a useful implementation check and the only universally applicable method to test different RNGs, it is an essential part of obtaining a cryptographic RNG certified by standards bodies—for example, the National Institute of Standards and Technology (NIST) or the Bundesamt für Sicherheit in der Informationstechnik (BSI). This certification process ensures that the RNG has been constructed and tested following best practices. For nearly all cryptographic applications, companies typically regard certification from a standards body as a prerequisite for RNG usage. The NIST and BSI standards require both the detailed modelling of the underlying physical process and the numerical testing of the RNG’s output statistics in order for a hardware RNG to be compliant. This has led to the development of statistical test suites, the best known being NIST’s Rukhin et al. (2001) and the Dieharder Marsaglia (2008); Brown et al. (2018) suite, but there are also others, e.g., L’ecuyer and Simard (2007); Walker ; Doty-Humphrey (2018). Together, these test suites enable the comprehensive analysis of a wide array of statistical characteristics. Despite their usefulness, they are often complicated to use and their outputs are challenging to analyse.
@@
To address this, we design a tunable statistical testing environment ($\mathsf{STE}$) by combining existing suites, allowing for adjustable intensity levels that balance the computational cost with effectiveness in detecting statistical bias, enabling more rigorous RNG testing than standard requirements. We make our testing environment openly available at https://github.com/CQCL/random_test (accessed on 28 November 2024). Using the $\mathsf{STE}$’s most intense setting, we test the output of three different RNGs representative of those used across commercial applications, allowing us to benchmark them against each other. We also provide a framework to analyse the overall numerical test results generated by the $\mathsf{STE}$, which is not obvious otherwise.
@@
We then present a range of post-processing methods, utilising various randomness extractors, to enhance the quality of each RNG’s output by, for instance, removing bias or dependencies between bits. Our set-up is illustrated in Figure 1. Each extraction method is based on distinct assumptions, which can be compared against each other. Examples of these assumptions include specific structures in the RNG’s output process, such as assuming that each output bit is generated in an identically and independently distributed (i.i.d.) manner or the availability of a short, pre-existing (near-)perfect random bit string as a resource. To evaluate the effectiveness of each method, we implement them using the Cryptomite software library Foreman et al. (2024) and apply statistical testing through our $\mathsf{STE}$ to analyse the impact of these techniques. These assumptions must be practically justified, and statistical testing is employed to determine whether each post-processing method succeeds in producing statistically sound results.
@@
Figure 1: This figure illustrates our implementation set-up. The black box represents one of the initial RNGs that we test, and the dashed box denotes the new—in principle, improved—RNG with additional post-processing applied.
1.1 Related Work
@@
The statistical testing of RNGs has a long history, dating back to the implementation of the Diehard CD-ROM tests in the 1990s Marsaglia (2008). Since then, two main research directions have emerged, which are both relevant to our work. First, researchers developed other test suites, such as the NIST Statistical Test Suite Rukhin et al. (2001), the TestU01 suite L’ecuyer and Simard (2007), ENT Walker , and PractRand Doty-Humphrey (2018), all of which we utilise for this work. Second, the empirical testing of specific RNGs was performed, such as Soto (1999); Tsvetkov (2011), which analysed and compared the results of statistical tests on a variety of pseudo-RNGs (PRNGs). Other works have considered the empirical testing of so-called true-RNGs (TRNGs)—for example, Hamburg et al. (2012), which tests the statistical properties of the entropy source in Intel’s Ivy Bridge TRNG, and Jun and Kocher (1999); Tsoi et al. (2007); Zhang et al. (2017); Williams et al. (2010); Sun and Lo (2018); Cho et al. (2020), which develop, implement, and statistically test different TRNGs. Recently, this has been extended to quantum RNGs (QRNGs)— for example, Xu et al. (2019); Ó Dúill et al. (2023); Jacak et al. (2021); Keshavarzian et al. (2023); Hurley-Smith and Hernandez-Castro (2020), in which the authors extensively statistical test an ID Quantique QRNG Quantique (2004). Other works give a universally applicable RNG statistical testing framework, such as Crocetti et al. (2023); Seyhan and Akleylek (2022).
@@
Randomness extraction also has a rich body of literature; see Shaltiel (2011) for an introduction. In cryptographic randomness standards, e.g., in NIST’s SP 800 Rukhin et al. (2001), so-called conditioners are standardised, whose role is similar to that of randomness extractors (randomness extractors can be understood as conditioners that have information-theoretic security, i.e., they do not rely on computational assumptions on the adversary). These conditioners are the only post-processing that has been vetted for use by governing bodies and are the most commonly used as a consequence. To the best of our knowledge, our work represents the first attempt at comparing the effects (from a statistical perspective) of different post-processing methods.
1.2 Summary of Results
In summary, our main results and observations are as follows.
@@ 1.2 Summary of Results
  • We perform intense statistical testing of three different RNGs: the 32-bit Linear Feedback Shift Register (LFSR) PRNG, Intel’s RDSEED TRNG, and IDQuantique’s ‘Quantis’ QRNG. We show the failure of two of them and provide evidence that one behaves well from a statistical perspective, extending and confirming the results of Hurley-Smith and Hernandez-Castro (2020); Hamburg et al. (2012).
@@
  • We present and implement a variety of post-processing methods, in the form of randomness extractors, to improve the output quality from each of the three RNGs. The post-processing methods form a set of four levels, where each level requires increasingly more sophisticated implementation: deterministic (level 1), seeded (level 2), two-source (level 3), and physical (level 4) extractors. Our contribution goes significantly further than the study and comparison of different types of extractors in Kwok et al. (2011); Ma et al. (2013), which focus only on deterministic or seeded extractors, respectively.
  • We present and experimentally demonstrate the implementation of a physical extractor (level 4) using a high-fidelity quantum computer. This allows us to execute a complex quantum protocol that has no classical equivalent.
@@
  • We intensively test the statistical effect that each level of post-processing has on the output of the different RNGs. Our main observations are as follows.
    • RNGs that fail statistical testing without post-processing continue to fail when simple post-processing methods (level 1) are applied, although some improvement is observed. Notably, one of these failing post-processed RNGs implements the self-shrinking generator Meier and Staffelbach (1994), which is studied for cryptographic applications.
@@
    • All our level 2, 3, and 4 implementations successfully post-process the output of the three RNGs used from a statistical perspective, passing the statistical testing.
@@
    • Low-entropy sources, such as the post-processed 32-bit LFSR, can pass rigorous statistical testing when suitable post-processing is applied. While this result aligns with the existence of cryptographically secure PRNGs, the low quality of the PRNG used highlights the limitations of statistical testing alone, i.e., without a precise model and justification for the unpredictability of the underlying physical process.
  • We make publicly available our $\mathsf{STE}$, a powerful, flexible, and easy to use statistical testing environment, together with suggested settings, which provide a valuable trade-off between the intensity and computation time, and a framework to analyse the cumulative test results.
2 Tools and Definitions
@@
As discussed in the Introduction, randomness is different when considering the perspective of a user or that of an adversary. In this work, we study the statistical properties of different RNGs’ outputs directly, without considering any additional information that a potential adversary might obtain, such as a detailed model of the entropy source or information from invasive or non-invasive attacks. We then examine, in the same manner, the effects of various post-processing methods applied to the RNG’s output. We consider the specific case of bits, i.e., the RNG outputs a bit string of length $n$, denoted $X\in\{0,1\}^{n}$, although one can also study RNGs whose output alphabet is larger. In some cases, we refer to sizes in bytes, where one byte equals eight bits. We denote the random variable produced by an RNG as $X$ and its specific realisation as $x$ (i.e., $X=x$). The set-up is illustrated in Figure 2.
@@
Figure 2: An illustration of the set-up that we consider. An RNG generates a bit string $X=x$ of length $n$. In this work, we first study the statistical properties of the realisation $x$ of the (random variable) $X$. Then, we analyse the effects of different post-processing methods applied to it.
@@
The amount of randomness that a random variable $X$ has is captured by its min-entropy $\mathrm{H}_{\infty}(X)$. {Definition}[Min-entropy] The min-entropy, $k$, of a random variable, $X\in\{0,1\}^{n}$, is defined as
@@
This can be interpreted as the minimum amount of randomness, in bits, that a variable $X$ has when there is no side information available. The min-entropy rate of a random variable $X\in\{0,1\}^{n}$ is $\alpha=\mathrm{H}_{\infty}(X)/n$. This can be interpreted as the minimum amount of randomness that $X$ has per bit, on average. Since RNGs output sequentially, we generalise this definition to consider the min-entropy of the current random variable conditioned on all previously produced random variables. This is known as block min-entropy.
{Definition}
[Block min-entropy] A set of random variables $X_{i}\in\{0,1\}^{n_{i}}$ for $i\in\mathbb{N}$ is said to have block min-entropy $k_{i}$, if
$\mathrm{H}_{\infty}(X_{i}|X_{0},X_{1},...,X_{i-1})=-\log_{2}\max\limits_{x\in\{0,1\}^{n_{i}}}\Pr(X_{i}=x|X_{0},X_{1},...,X_{i-1})\geq k_{i},\hskip 19.91684pt\forall i\ .$ (2)
This can be interpreted as the minimum number of random bits that a variable $X_{i}$ has when conditioned on all previous random variables, indexed by $10,\ldots,i-$ 1.
{Definition}
[Statistical distance] The statistical distance, $\Delta$, between two random variables, $X,Z\in\{0,1\}^{n}$, is defined as
This is a measure of how close to one another, or indistinguishable from one another, two random variables are.
{Definition}
[$\epsilon$-perfect randomness] A random variable $X$ on $\{0,1\}^{n}$ is said to be $\epsilon$-perfectly random if
$\Delta(X,\mathrm{U}_{n})\leq\epsilon\ ,$ (4)
where $\mathrm{U}_{n}$ is the uniform variable on $\{0,1\}^{n}$, i.e., $\Pr(\mathrm{U}_{n}=u)=\frac{1}{2^{n}}$ for all $u\in\{0,1\}^{n}$. This definition is equivalent to saying that the variable $X$ is distinguishable from a uniform distribution with a distinguishing advantage of at most $\epsilon$, i.e., a distinguisher can guess that $X$ is not uniform with a success probability of at most $\frac{1}{2}+\epsilon$. When $\epsilon=0$, the random variable is said to be perfectly random. This definition is universally composable Canetti (2001), i.e., $X$ can be used safely in other applications.
3 Statistical Testing
Statistical test suites are collections of algorithms that analyse the numerical properties of a set of random numbers to determine whether there is evidence to reject the possibility that they are uniformly distributed. If there is sufficient evidence to reject this possibility, a statistical test is said to be failed and the RNG output can be distinguished from the uniform distribution at some confidence level. The hypothesis that a random variable is uniformly distributed is known as the null hypothesis $\mathsf{H_{0}}$. For an RNG producing a random variable $X\in\{0,1\}^{n}$, the null hypothesis is $\mathsf{H_{0}}:\Delta(X,\mathrm{U}_{n})=0$. If the null hypothesis is rejected, then the alternative hypothesis $\mathsf{H_{1}}:\Delta(X,\mathrm{U}_{n})>0$ is accepted.
However, a random variable cannot be tested directly; only its realisation can—i.e., the bit string $x\in\{0,1\}^{n}$ produced by the random variable $X$. To assess whether to accept or reject the null hypothesis, a statistical test calculates a specific measure of $x$ (e.g., its mean), known as the test statistic $t$, and analyses how likely this test statistic is to be observed, assuming that the underlying random variable is uniform. Test statistics calculated from realisations of a uniform distribution are normally distributed, so one can calculate how likely observing certain ranges of the test statistic is by using concentration inequalities. More precisely, this likelihood is captured by a probability known as the p-value, which is defined as follows. {Definition}[p-value] Given an observed test statistic $t$ obtained by calculating a measure from the realisation of a random variable $X=x\in\{0,1\}^{n}$ and $T$, the (normally distributed) variable associated with all the possible measure values, the p-value $\mathsf{p}\in[0,1]$ is defined as
$\mathsf{p}=\Pr(T\leq t|\Delta(X,\mathrm{U}_{n})=0)\ ,$ (5)
where $\mathrm{U}_{n}\in\{0,1\}^{n}$ is uniformly distributed.
A range of p-values is defined that provides a threshold at which the null hypothesis is rejected, i.e., when the test is deemed to fail. If a test ensures that there is, at most, a $1\%$ chance that it incorrectly rejects that the RNG is producing uniform random numbers (known as the type 1 error—when a statistical test incorrectly rejects a true null hypothesis), then it would, for example, conclude failure if $\mathsf{p}\notin[0.01,1]$. This threshold for failure is on one tail only, so it only fails test statistics that are sufficiently biased away from the expected value in one direction. More generally, tests are two-tailed and conclude failure if the observed p-values are outside of a sufficiently large interval—for example, if $\mathsf{p}\notin[0.005,0.995]$.
The failure of numerous statistical tests is a strong indicator that an RNG is not producing (near-)perfect random numbers, as its output can be distinguished from the uniform distribution with a high probability. For example, if all statistical tests performed on the RNG are independent, the probability that the null hypothesis is accepted given that the alternative hypothesis is true (known as the type 2 error) is $\mathsf{p_{type2}}=\mathsf{p^{test1}_{type2}}\cdot\mathsf{p^{test2}_{type2}}\cdot\ldots\cdot\mathsf{p^{testn}_{type2}}$, where $n$ is the number of tests performed.
We now describe several existing statistical test suites used in this work. It is important to note that, while each suite contains multiple tests, many test outcomes are correlated. For instance, a source with an unusually high number of ones may fail both the monobit and poker tests. Similarly, different test suites often include the same tests with slight parameter variations, meaning that failure in one suite is likely to result in failure in another. For example, both NIST and Dieharder include a runs test.
3.1 Existing Test Suites
3.1.1 NIST Statistical Test Suite
The NIST statistical test suite (SP 800-22) Rukhin et al. (2001) is the best known. It includes 15 tests, some with multiple sub-tests, and passing certain tests is required for RNG certification by organisations such as NIST and BSI. During testing, a randomness file is split into sub-strings, with each sub-string tested individually. Users can specify the number of sub-strings and the total bit string size to analyse, although the guide recommends using 100 sub-strings of $10^{6}$ bits, requiring at least $10^{8}$ bits, or 12.5 MB, for testing. For each test, an analysis is performed on each sub-string, and the suite provides two results: (1) the p-value for a statistical test on the uniformity of the distribution of results across sub-strings and (2) the number of sub-strings that pass each test. Both results are assessed at the 1% significance level. We note that the individual p-values for each sub-string in each test are not accessible to the user.
3.1.2 Diehard(er) Statistical Test Suite
The Dieharder statistical test suite includes the 18 original Diehard tests along with additional tests, including some from the NIST suite. Like the NIST suite, it is widely used by RNG certification bodies. A failure is determined when $\mathsf{p}\notin[0.0005,1-0.0005]$, and a test is considered ’weak’ if $\mathsf{p}\in[0.0005,0.005]\cup[1-0.0005,1-0.005]$. This higher tolerance for poor test statistics means that a flawed RNG may occasionally pass Dieharder, but failure is a strong indicator of non-uniformity. The Dieharder tests require a large quantity of random numbers to prevent the re-use of input data, which can lead to inaccurate results. We recommend using at least 1 gigabyte (GB) of random numbers for testing. For smaller file sizes, the test parameters can be adjusted to avoid these issues. In our testing environment, we use the default parameters for each test.
3.1.3 TestU01 Statistical Test Suite
TestU01 is a C-based software library for conducting RNG statistical testing with pre-compiled test batteries. These batteries vary significantly in the number of tests and the amount of randomness required. For details on the specific tests included in each battery, see L’ecuyer and Simard (2007). Test p-values are reported if $\mathsf{p}\notin[0.001,0.999]$, which we use as our failure criterion. In our testing, we use the Alphabit, Rabbit, and SmallCrush batteries from TestU01. To run these tests, the input files must contain at least $2^{25}$ random bytes (approximately 35 MB). We omit the Crush and BigCrush batteries due to their long runtimes and large randomness requirements, although they can be run within our statistical testing environment.
3.1.4 ENT Statistical Test Suite
The ENT test suite is a small but efficient set of six statistical tests. It has been used to demonstrate bias in a commercial quantum RNG by consistently failing the $\chi^{2}$ test Hurley-Smith and Hernandez-Castro (2017) (we replicate these results with an independently acquired RNG; see Table 3. ENT outputs test statistics without providing a pass/fail threshold, so we assess failure based on the criteria in Table 3 of Ortiz-Martin et al. (2018). Although there is no specific guidance on the required input sizes, we found that the tests produced unreliable results with inputs smaller than $0.5$ GB.
3.1.5 PractRand Statistical Test Suite
PractRand is a C++  library of statistical tests designed for practicality—they are efficient, user-friendly, and capable of detecting significant biases in RNGs. According to its documentation, it runs faster than most test suites (which we confirm; see Table 2), offers unique interfacing, has no theoretical maximum input length, and includes some original tests. PractRand tests input files based on size, examining subsets of $2^{24+x}$ bytes for $x\in\mathbb{N}$, with more tests performed as $x$ increases. In our testing, we limit the maximum test size to $2^{32}$ bytes (approximately 4.3 GB). For full details and comparisons with other test suites, see Doty-Humphrey (2018). PractRand uses various p-value ranges, including “unusual”, “mildly suspicious”, “suspicious”, “suspect”, and “fail”. A failure occurs when $\mathsf{p}\notin[10^{-11},1-10^{-11}]$.
3.2 Our Statistical Testing Environment
The interfacing code for our $\mathsf{STE}$ can be downloaded at https://github.com/CQCL/random_test (accessed on 28 November 2024). We offer three testing modes, Light, Recommended, and All, which can be executed using the commands run_light, run_recommended, and run_all, respectively. The NIST statistical test suite is not included in these commands due to its need for user prompts, but it can be run separately within the environment. We believe that the Recommended setting strikes a good balance between the computational (and environmental) cost and rigor, exceeding the standard testing required by certification bodies. All results in this work can be replicated using the $\mathsf{STE}$ or by downloading, configuring, and running the relevant statistical test suites independently.
3.2.1 Suggested Settings
We now propose the recommended settings for statistical testing using our $\mathsf{STE}$, based on insights gained during this research. The runtimes are averaged over 10 executions using a 10 Gbit file, except for the NIST suite, where a 100 Mbit file is tested, in accordance with the user guidelines. All testing was conducted on a Dell Precision 7540 laptop with 16 GB of RAM and a 2.3 GHz Intel I9-9880H processor, running the Ubuntu 20.04 operating system. All the test runtimes are given in Tables 1 and 2.
Table 1: This table details our settings for light, recommended, and all statistical testing using the code provided. A ‘Y’ in a specific column indicates that the associated test suite of this column is included in the setting.
{adjustwidth}